OpenStudy (anonymous):

what is the equation of a line that passes through (-2,3) and is parallel to 2x+3y=6? what is the equation of a line that passes through (-2,3) and is parallel to 2x+3y=6? @Mathematics

OpenStudy (amistre64):

parallel lines have the same equations, only differ at most by a constant

OpenStudy (amistre64):

2x+3y=6 2x+3y=C -2 3 --------- -4 + 9 = 5 2x+3y=5